OpenStudy (anonymous):

How do I solve 4x + 3y = -1 3x +y = 3 using subsitution

OpenStudy (anonymous):

4x + 3y = -1 3x +y = 3 1st solve for x or y in the first equation i'll solve for x in first one x = -(3/4)*y-1/4 2nd plug that into second eq and solve for the only variable. In this case it is y 3(-(3/4)*y-1/4)+y= 3 y=-3 :) 3rd since now we have y=-3 we can plug that into first eq or second eq and just solve for the remaining variable which is x. i'll pick the second eq 3x +y = 3 3x+(-3)=3 x=2 so solution is x=2 y=-3

OpenStudy (anonymous):

Thank you soo much :) I appreciate it!!!

mathslover (mathslover):

@heather96 Welcome to openstudy